A rectangular flower bed with the width 12 ft. and length 15 ft. is going to have a walkway of uniform width on all sides. If area of the walkway is 90 sq. ft., how wide is the walkway?
Answer by nyc_function(2741)

Find x.
Here's another way to solve for x.
Let x = width of walkway
(12 - 2x)(15 - 2x) = total area or 180
180 - 54x + 4x^2 = 180
180 - 180 - 54x + 4x^2 = 0
-54x + 4x^2 = 0
x = 27/2